# Geometry induced net spin polarization of $d$-wave altermagnets

### v1

@@ Abstract
Altermagnets exhibit spin-split electronic band structures despite having zero net magnetization, making them attractive for field-free spintronic applications. In this work, we show that a finite rectangular altermagnetic sample can acquire a net spin polarization purely due to its geometry. This effect arises from the interplay between the anisotropic, spin-resolved Fermi contours of an altermagnet and the discrete sampling of momentum space imposed by unequal sample dimensions. By explicitly counting occupied states, we demonstrate that rectangular samples with $L_{x}\neq L_{y}$ host a finite spin polarization, which vanishes in the symmetric limit $L_{x}=L_{y}$ and in the thermodynamic limit. We further show that this geometry-induced spin polarization can be directly probed in transport measurements. In the tunneling regime, the ratio of spin to charge conductance exhibits characteristic square-like patterns as a function of sample dimensions, faithfully reflecting the underlying spin polarization. In addition, transport across ferromagnet–altermagnet–ferromagnet junctions reveals an asymmetric magnetoresistance with respect to reversal of the Zeeman field, providing an independent transport signature of the finite spin polarization. Our results establish geometry as an effective control parameter for spin polarization in altermagnets and suggest a viable route for exploiting finite-size effects in mesoscopic altermagnetic spintronic devices.
@@ I Introduction
Beyond these effects, the anisotropic nature of the spin-resolved band structures in AMs provides an additional and largely unexplored route to controlling spin polarization. In particular, when the electronic spectrum is sampled discretely, as in finite-sized systems, the anisotropy of the Fermi contours can lead to an unequal population of the two spin species. This opens up the possibility of inducing a net spin polarization purely through sample geometry, without invoking external magnetic fields or intrinsic magnetization. In this work, we analyze this geometry-induced spin polarization in finite altermagnetic samples and demonstrate how it manifests itself in transport measurements. We further propose experimentally accessible schemes to detect this spin polarization, highlighting the role of finite-size and shape effects as effective control parameters in altermagnet-based spintronic devices.
@@ II Spin polarization of a two-dimensional rectangular AM
$H$ $=$ $-ta^{2}(\partial_{x}^{2}+\partial_{y}^{2})\sigma_{0}-t_{J}a^{2}(\partial_{x}^{2}-\partial_{y}^{2})\sigma_{z},$ (1)
@@
Figure 1: Normalized spin polarization $(N_{\uparrow}-N_{\downarrow})/(N_{\uparrow}+N_{\downarrow})$ versus $(L_{x},L_{y})$ for $t_{J}=0.8t$ and $E_{F}=0.6t$. The samples with $L_{x}\neq L_{y}$ exhibit nonzero spin polarization.
@@ III Transport probe of geometry-induced spin polarization
$H$ $=$ $\begin{cases}-t_{n}a^{2}(\partial_{x}^{2}+\partial_{y}^{2}+\partial_{z}^{2})\sigma_{0},~~~{\rm for~~~}|z|>d/2,\\ ~~\\ -ta^{2}(\partial_{x}^{2}+\partial_{y}^{2}+\partial_{z}^{2})\sigma_{0}-t_{J}a^{2}(\partial_{x}^{2}-\partial_{y}^{2})\sigma_{z}\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~{\rm for~~}|z|<d/2,\end{cases}$ (3)
@@
in addition to the continuity of $\psi$. For all $z$, the system is confined to $0<x<L_{x}$ and $0<y<L_{y}$. We impose periodic boundary conditions along the $x$ and $y$ directions, so that $k_{x}$ and $k_{y}$ remain good quantum numbers. Along the $z$ direction, the AM slab is subject to open boundary conditions. As a consequence, the lowest allowed transverse mode acquires a finite energy shift $ta^{2}\pi^{2}/d^{2}$ compared to the strictly two-dimensional case, where $d$ denotes the thickness of the AM slab.
Figure 3: The ratio of spin conductance to charge conductance (both in dimensionless units) across AM versus the system size $(L_{x},L_{y})$ for $t_{n}=0.2t$, $t_{J}=0.8t$, and $d=5a$.
The transport properties are evaluated using the Landauer formalism, with the conductance calculated independently in the two spin sectors. The sum of the two contributions defines the charge conductance $G$, while their difference yields the spin conductance $G^{s}$. The ratio $G^{s}/G$ therefore provides a direct and physically transparent measure of the spin polarization in transport. In Fig. 3, we plot $G^{s}/G$ versus $(L_{x},L_{y})$ for $q_{0}=10$, $E_{F}=t$, with $t_{n}=0.2t$, $t_{J}=0.8t$, and $d=5a$. The hopping parameter $t_{n}$ in the normal-metal leads is chosen such that their Fermi surface fully encloses that of the AM, ensuring that all relevant AM bands participate in transport.
As evident from Fig. 3, the ratio $G^{s}/G$ exhibits a square-like pattern whose characteristic size closely matches that obtained from the state-counting analysis in Fig. 1. The sharp boundaries between different regions correspond to abrupt changes in the net spin polarization of AM and are faithfully reproduced by the transport calculations. We emphasize that this direct correspondence between the spin polarization and transport signatures emerges only in the tunneling limit.
@@
Figure 4: Magnetoresistance of the FM-AM-FM setup. (a) MR versus Zeeman energy $b$ for different system sizes. (b) MR versus $L_{0}$ for system with size $L_{0}\times 5L_{0}$ for $b=0.5t$. Other parameters are same as in Fig. 3.
We compute the conductance $G(b)$ as a function of the Zeeman energy $b$ and characterize the transport response through the magnetoresistance (MR), defined as the relative change in resistance between finite and zero Zeeman fields. In Fig. 4(a), we plot the MR as a function of $b$ for several choices of the aspect ratio $L_{x}/L_{y}$, while keeping the cross-sectional area approximately fixed. For rectangular samples with $L_{x}\neq L_{y}$, the MR is clearly asymmetric under $b\to-b$, reflecting the presence of a nonzero net spin polarization in the AM. In contrast, for $L_{x}=L_{y}$, the MR is symmetric in $b$, consistent with the absence of net spin polarization.
In Fig. 4(b), we also show the dependence of MR on the system size for fixed values of $b$ and the aspect ratio $L_{x}/L_{y}$. The MR decreases with increasing system size and approaches zero in the large-size limit. This behavior can be traced back to the scaling of the net spin polarization in the AM: although the anisotropic band structure favors unequal population of the two spin species in a finite rectangular sample, the total number of available states for each spin scales with the product $L_{x}L_{y}$. Consequently, the spin polarization per unit area vanishes in the thermodynamic limit, leading to a suppression of the MR for large systems.
@@ IV Summary and conclusions
We demonstrated that this geometry-induced spin polarization can be directly detected in transport measurements. In the tunneling regime, the spin polarization inferred from state counting is faithfully reflected in the ratio of spin to charge conductance, $G^{s}/G$, which exhibits characteristic square-like patterns as a function of the sample dimensions. These features establish a direct correspondence between the underlying band-structure anisotropy and experimentally accessible transport observables.
Furthermore, we investigated transport across FM–AM–FM junctions and showed that the presence of a net spin polarization in the AM manifests itself as an asymmetry in the magnetoresistance under reversal of the Zeeman field in the ferromagnetic leads. This asymmetry disappears for square samples, consistent with the vanishing of the geometry-induced spin polarization when $L_{x}=L_{y}$. We also found that the magnetoresistance decreases with increasing system size, reflecting the fact that the spin polarization per unit area vanishes in the thermodynamic limit.
@@ Acknowledgements.
The author thanks Science and Engineering Research Board (now Anusandhan National Research Foundation) Core Research grant (CRG/2022/004311) and University of Hyderabad for financial support.

### v2

@@ Abstract
Altermagnets exhibit spin-split electronic bandstructures despite having zero net magnetization, making them attractive for field-free spintronic applications. In this work, we show that a finite rectangular altermagnetic sample can acquire a net spin polarization purely due to its geometry. This effect arises from the interplay between the anisotropic, spin-resolved Fermi contours of an altermagnet and the discrete sampling of momentum space imposed by unequal sample dimensions. By explicitly counting occupied states, we demonstrate that rectangular samples with $L_{x}\neq L_{y}$ host a finite spin polarization, which vanishes in the symmetric limit $L_{x}=L_{y}$ and in the thermodynamic limit. We further show that this geometry-induced spin polarization can be directly probed in transport measurements. In the tunneling regime, the charge and the spin conductances exhibit characteristic patterns as a function of sample dimensions, faithfully reflecting the underlying spin polarization. In addition, transport across ferromagnet–altermagnet–ferromagnet junctions reveals an asymmetric magnetoresistance with respect to reversal of the Zeeman field, providing an independent transport signature of the finite spin polarization. Our results establish geometry as an effective control parameter for spin polarization in altermagnets and suggest a viable route for exploiting finite-size effects in mesoscopic altermagnetic spintronic devices.
@@ I Introduction
Beyond these effects, the anisotropic nature of the spin-resolved bandstructures in altermagnets provides an additional and largely unexplored route to controlling spin polarization. In particular, when the electronic spectrum is sampled discretely, as in finite-sized systems, the anisotropy of the Fermi contours can lead to an unequal population of the two spin species. This opens up the possibility of inducing a net spin polarization purely through sample geometry, without invoking external magnetic fields or intrinsic magnetization. Related finite-size effects have been identified in altermagnetic strips, where a net local magnetization develops near the edges [4]. In contrast, we focus on a geometry-induced spin polarization that arises from the global shape and aspect ratio of the sample and manifests itself in transport measurements. Specifically, we analyze this effect in finite altermagnetic samples and propose experimentally accessible schemes to detect it, highlighting finite-size and shape effects as effective control parameters in altermagnet-based spintronic devices.
@@ II Spin polarization of a two-dimensional rectangular AM
$H$ $=$ $-ta^{2}(\partial_{x}^{2}+\partial_{y}^{2})\sigma_{0}+t_{J}a^{2}(\partial_{x}^{2}-\partial_{y}^{2})\sigma_{z},$ (1)
@@
Figure 1: Normalized spin polarization $(N_{\uparrow}-N_{\downarrow})/(N_{\uparrow}+N_{\downarrow})$ versus $(L_{x},L_{y})$ for $t_{J}=0.8t$ and $E_{F}=0.6t$ with periodic boundary conditions. The samples with $L_{x}\neq L_{y}$ exhibit nonzero spin polarization.
@@
When periodic boundary conditions are imposed, the allowed momenta are quantized as $k_{j}=n_{j}\pi/L_{j}$ for $j=x,y$, where $n_{j}$ takes integer values, including zero. The condition $E(\vec{k})<E_{F}$ then translates into $|n_{j}|<k_{F,j}L_{j}/\pi$ for modes with $n_{j^{\prime}}=0$, where $j^{\prime}=y,x$ for $j=x,y$, and the spin-dependent Fermi wavevectors are given by $k_{F,x}=\sqrt{E_{F}/[(t-s_{\sigma}t_{J})a^{2}]}$ and $k_{F,y}=\sqrt{E_{F}/[(t+s_{\sigma}t_{J})a^{2}]}$. Consequently, the square-like features observed in Fig. 1 originate from modes with either $n_{x}=0$ or $n_{y}=0$. This behavior changes qualitatively when open boundary conditions are imposed, for which $n_{j}$ is restricted to positive integers and the $n_{j}=0$ modes are excluded. As a result, the square-like pattern disappears in the normalized spin polarization plotted as a function of sample dimensions, as shown in Fig. 2. This distinction is particularly relevant for experiments, since periodic boundary conditions are an idealization and realistic finite samples are governed by open boundary conditions, for which the square-like features are absent.
Figure 2: (a) Normalized spin polarization $(N_{\uparrow}-N_{\downarrow})/(N_{\uparrow}+N_{\downarrow})$ (b) $(N_{\uparrow}+N_{\downarrow})$ (c) $(N_{\uparrow}-N_{\downarrow})$, versus the sample dimensions $(L_{x},L_{y})$. The samples with $L_{x}\neq L_{y}$ exhibit nonzero spin polarization.
@@ III Transport probe of geometry-induced spin polarization
$H$ $=$ $\begin{cases}-t_{n}a^{2}(\partial_{x}^{2}+\partial_{y}^{2}+\partial_{z}^{2})\sigma_{0},~~~{\rm for~~~}|z|>d/2,\\ ~~\\ -ta^{2}(\partial_{x}^{2}+\partial_{y}^{2}+\partial_{z}^{2})\sigma_{0}+t_{J}a^{2}(\partial_{x}^{2}-\partial_{y}^{2})\sigma_{z}\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~{\rm for~~}|z|<d/2,\end{cases}$ (3)
@@
in addition to the continuity of $\psi$. For all $z$, the system is confined to $0<x<L_{x}$ and $0<y<L_{y}$. We impose open boundary conditions along the $x$ and $y$ directions, as these provide a more realistic description of a finite sample. Along the $z$ direction, the AM slab is subject to open boundary conditions. As a consequence, the lowest allowed transverse mode acquires a finite energy shift $ta^{2}\pi^{2}/d^{2}$ compared to the strictly two-dimensional case, where $d$ denotes the thickness of the AM slab.
Figure 4: (a) Charge conductance in units of $e^{2}/h$ and (b) spin conductance in units of $e/2$ across AM versus the system size $(L_{x},L_{y})$ for $t_{n}=0.2t$, $t_{J}=0.8t$, and $d=5a$ for the system with open boundary conditions along $x$ and $y$.
The transport properties are evaluated using the Landauer formalism, with the conductance calculated independently in the two spin sectors. The sum of the two contributions defines the charge conductance $G$, while their difference yields the spin conductance $G^{s}$. In Fig. 4, we plot (a) $G$ and (b) $G_{s}$ versus $(L_{x},L_{y})$ for $q_{0}=10$, $E_{F}=t$, with $t_{n}=0.2t$, $t_{J}=0.8t$, and $d=5a$. The hopping parameter $t_{n}$ in the normal-metal leads is chosen such that their Fermi surface fully encloses that of the AM, ensuring that all relevant AM bands participate in transport.
As evident from Fig. 4, both the charge conductance $G$ and the spin conductance $G_{s}$ exhibit extrema along contours that coincide with the transitions in the number of occupied states shown in Fig. 2. Along these contours, the energy of a state in the isolated AM sample matches the energy of electrons incident from the leads, giving rise to transmission resonances. We emphasize that this direct correspondence between spin polarization and transport signatures emerges only in the tunneling limit.
@@
Figure 5: Magnetoresistance (MR) of the FM–AM–FM junction. (a) MR as a function of the Zeeman energy $b$ for different system sizes. Each curve corresponds to a distinct choice of $(L_{x},L_{y})/a$, as indicated in the legend. (b) MR versus $b$ for systems with the same aspect ratio, $(L_{x},L_{y})=n(4,10)a$, for different values of $n$ shown in the legend. (c) The asymmetry measure $\mathrm{MR}(b)-\mathrm{MR}(-b)\,]/2b$ as a function of the characteristic system size $L_{0}$, for $b=0.1t$. It can be seen that in the limit of large system sizes the asymmetry measure tends to zero. All other parameters are the same as in Fig. 4.
We compute the conductance $G(b)$ as a function of the Zeeman energy $b$ and characterize the transport response via the magnetoresistance (MR), defined as $\mathrm{MR}=\delta R(b)/R(0)$, where $\delta R(b)=R(b)-R(0)$ and $R(b)=1/G(b)$ is the resistance at Zeeman energy $b$. In Fig. 5(a), we plot the MR as a function of $b$ for several choices of the aspect ratio $L_{x}/L_{y}$, while keeping the cross-sectional area fixed. For rectangular samples with $L_{x}\neq L_{y}$, the MR is clearly asymmetric under $b\to-b$, reflecting the presence of a nonzero net spin polarization in the altermagnet. In contrast, for $L_{x}=L_{y}$, the MR is symmetric in $b$, consistent with the absence of net spin polarization.
A symmetric MR–$b$ curve indicates the absence of net spin polarization, whereas an asymmetric dependence signals a finite spin polarization. In Fig. 4(b), we plot the MR as a function of $b$ for several system sizes sharing the same aspect ratio. As the system size increases, the slope of the MR curve at $b=0$ is progressively reduced. We quantify this asymmetry using the slope at $b=0$, which may be approximated by $\mathrm{MR}(b)-\mathrm{MR}(-b)\,]/2b$ for small $b$. Figure 4(c) shows this asymmetry measure as a function of system size for a fixed aspect ratio. In the limit of large system sizes, the asymmetry tends to zero, indicating that the net spin polarization in altermagnets is predominantly a finite-size effect.
@@ IV Summary and conclusions
We demonstrated that this geometry-induced spin polarization can be directly detected in transport measurements. In the tunneling regime, the spin polarization inferred from state counting is faithfully reflected in the charge and spin conductances which exhibit characteristic patterns as a function of the sample dimensions. These features establish a direct correspondence between the underlying band-structure anisotropy and experimentally accessible transport observables.
Furthermore, we investigated transport across FM–AM–FM junctions and showed that the presence of a net spin polarization in the AM manifests itself as an asymmetry in the magnetoresistance under reversal of the Zeeman field in the ferromagnetic leads. This asymmetry disappears for square samples, consistent with the vanishing of the geometry-induced spin polarization when $L_{x}=L_{y}$. We also found that the asymmetry decreases with increasing system size, reflecting the fact that the spin polarization per unit area vanishes in the thermodynamic limit.
@@ Acknowledgements.
The author thanks Jacob Linder and Pavlo Sukhachov for useful discussions. The author thanks Science and Engineering Research Board (now Anusandhan National Research Foundation) Core Research grant (CRG/2022/004311) and University of Hyderabad for financial support.